If 18 people out of 8,400 people responded to a sales offer, how many of the 8,400 people read the offer but did not respond? I am seeking a probability statistic such as used in a "coin toss", not subtraction (i.e. not 8,382 as the answer). Thank you
Answer #1

i think that you are asking about the proportion of people who read the offer but not respond which is equal to - 1-prob. of people who respond.

now prob. of people who respond =no. of people who respond/total no. of people =18/8400=0.0021

therefore required probability or proportion of people who not respond =1-0.0021 = 0.9979